OpenStudy (anonymous):

Frances has 4 square baking dishes. Each pan is 7 inches × 7 inches × 2 inches. Which expression will give her the total volume of all the pans?

OpenStudy (anonymous):

Volume is length•width• height

OpenStudy (aq99):

7*7*2=X X*4=Y Y is the total :)

OpenStudy (dayakar):

volume of each pan = L*B*H VOLUME OF 4 PANS = 4*L*B*H = 4*7*7*2 = 392CUBIC INCHES